# Fate of non-Hermitian free fermions with Wannier-Stark ladder

### v3

@@ I Introduction
The effective non-Hermitian Hamiltonian [26, 27], which evolves under continuous measurement without quantum jumps [28], could also serve as a viable setup for studying the MIET in nonunitary evolution scenarios [28, 29]. The most interesting case is the interplay between the non-Hermitian skin effect (NHSE) [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], characterized by a large number of eigenmodes being anomalously localized at one side of the open boundary due to nonreciprocal dissipation, and a free fermion system [52, 53]. There, weak measurements and post-selections encapsulated by the non-Hermicity [28], cause non-Hermitian free fermions to undergo a volume-to-area law entanglement scaling transition. Subsequently, [54] explored how the interaction between the NHSE in non-Hermitian free fermions and Anderson localization [55] leads to divergent scaling laws in log-to-volume-to-area phases under non-Hermitian disordered evolution. The similar phenomenon is also observed in the non-Hermitian Aubry-André-Harper chain [56], albeit with entanglement transitions that follow area-to-volume-to-area scaling laws. Meanwhile, [57] confirms transitions between typical volume-to-area scaling laws and area-to-log-to-volume scaling laws in two distinct non-Hermitian quasiperiodic systems. In addition, these similar physical issues, entanglement dynamical competition between localization and NHSE, also, are discussed in the context of disordered monitored free fermions [58]. Clarifying the intricate entanglement dynamics between the NHSE and localization significantly enhances our understanding of the MIET during nonunitary evolution.
@@
The rest of the paper is organized as follows. In Sec. II, we first introduce the effective Hamiltonian of non-Hermitian free fermions with WSL and the numerical methods employed. In Sec. III, we present the numerical results, identifies the physical quantities necessary for detecting entanglement phase diagram and transitions. We also give a brief discussion from the perspective of single-particle localization in Sec. IV. A conclusion and outlook is given in Sec. V. Additional numerical calculation data are included in the Appendix.
@@ II.1 Effective Hamiltonian of Non-Hermitian free fermions model with Wannier-Stark ladder
In this work, we consider a effective Hamiltonian [see Appendix A for the discussion of the effective non-Hermitian Hamiltonians] of one-dimensional spinless free fermions with Hatano-Nelson nearest neighbor (NN) hopping, subject to a WSL [see FIG. 1(a)]:
@@ III.2 Entanglement transitions and critical scaling properties
The inset of FIG. 2(c) fits the critical scaling exponent $S_{L/2}\sim L^{\beta}$, yielding $\beta\approx 0.922\pm 0.215$, consistent with the finite-size scaling results, $\zeta/\nu\approx 1.03(1)$ [77]. FIG. 2(d) shows the transition from volume-law to area-law behavior as a function of system size $L$ under a small gradient of WSL ($\Delta=0.001$), driven by increasing non-Hermitian strength $\gamma$ in the range $[-0.001,-0.1]$, due to the NHSE.
@@ IV Discussions
$|\psi_{j}|\propto\left\{\begin{array}[]{lc}e^{-(\tau(E)-g)(j-j_{0})}.,&j>j_{0},\\ e^{-(\tau(E)+g)(j_{0}-j)}.,&j<j_{0}.\\ \end{array}\right.$ (15)
@@ Acknowledgements.
We deeply appreciate Shan-Zhong Li, Shuo Liu and Ze-Chuan Liu for the valuable discussions and support, and H.-Z. Li thanks Wen Wang for their assistance on this work. H.-Z. Li also acknowledges Xue-Jia Yu for the guidance and support. Finally, we dedicate this paper to the vibrant and rapidly emerging Institute for Quantum Science and Technology (IQST) at Shanghai University. We gratefully acknowledges the National Natural Science Foundation of China (Grant No. 11874316), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Team in University (Grant No. IRT13093).

### v4

@@ I Introduction
The effective non-Hermitian Hamiltonian [26, 27], which evolves under continuous measurement without quantum jumps [28], could also serve as a viable setup for studying the MIET in nonunitary evolution scenarios [28, 29]. The most interesting case is the interplay between the non-Hermitian skin effect (NHSE) [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], characterized by a large number of eigenmodes being anomalously localized at one side of the open boundary due to nonreciprocal dissipation, and a free fermion system [52, 53]. There, weak measurements and post-selections encapsulated by the non-Hermicity [28], cause non-Hermitian free fermions to undergo a volume-to-area law entanglement scaling transition. Subsequently, [54] exploring how the interaction between the NHSE in non-Hermitian free fermions and Anderson localization [55] leads to divergent scaling laws in log-to-volume-to-area phases under non-Hermitian disordered evolution. The similar phenomenon was also observed in the non-Hermitian Aubry-André-Harper chain [56], albeit with entanglement transitions that follow area-to-volume-to-area scaling laws. Meanwhile, transitions between typical volume-to-area scaling laws and area-to-log-to-volume scaling laws in two distinct non-Hermitian quasiperiodic systems were confirmed  [57]. In addition, these similar physical issues, entanglement dynamical competition between localization and NHSE, are discussed in the context of disordered monitored free fermions [58]. Clarifying the intricate entanglement dynamics between the NHSE and localization significantly enhances our understanding of the MIET during nonunitary evolution.
@@
The rest of the paper is organized as follows. In Sec. II, we first introduce the effective Hamiltonian of non-Hermitian free fermions with WSL and the numerical methods employed. In Sec. III, we present the numerical results to identify the physical quantities necessary for detecting entanglement phase diagram and transitions. We also give a brief discussion from the perspective of single-particle localization in Sec. IV. Conclusion and outlooks are given in Sec. V. Additional numerical calculation data are included in the Appendix.
@@ II.1 Effective Hamiltonian of Non-Hermitian free fermions model with Wannier-Stark ladder
In this work, we consider an effective Hamiltonian [see Appendix A for the discussion of the effective non-Hermitian Hamiltonians] of one-dimensional spinless free fermions with Hatano-Nelson nearest neighbor (NN) hopping, subject to a WSL [see FIG. 1(a)]:
@@ III.2 Entanglement transitions and critical scaling properties
The inset of FIG. 2(c) fits the critical scaling exponent $S_{L/2}\sim L^{\beta}$, yielding $\beta\approx 0.922\pm 0.215$, consistent with the finite-size scaling results with $\zeta/\nu\approx 1.03(1)$ [77]. FIG. 2(d) shows the transition from volume-law to area-law behavior as a function of system size $L$ under a small gradient of WSL ($\Delta=0.001$), driven by increasing non-Hermitian strength $\gamma$ in the range $[-0.001,-0.1]$, due to the NHSE.
@@ IV Discussions
$|\psi_{j}|\propto\left\{\begin{array}[]{lc}e^{-(\tau(E)-g)(j-j_{0})},&j>j_{0},\\ e^{-(\tau(E)+g)(j_{0}-j)},&j<j_{0}.\\ \end{array}\right.$ (15)
@@ Acknowledgements.
We deeply appreciate Shan-Zhong Li, Shuo Liu and Ze-Chuan Liu for the valuable discussions and support, and H.-Z. Li thanks Wen Wang for her assistance on this work. H.-Z. Li also acknowledges Xue-Jia Yu for the guidance and support. Finally, we dedicate this paper to the vibrant and rapidly emerging Institute for Quantum Science and Technology (IQST) at Shanghai University. We gratefully acknowledges the National Natural Science Foundation of China (Grant No. 11874316), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Team in University (Grant No. IRT13093).